angle 1 and angle 2 are supplementary angles. m angle 1 is 4y + 7 and m angle 2 is 9y + 4. Find m angle 2.
Answer by cleomenius(959)

4y + 7 + 9y + 4 = 180
13y + 11 = 180
13y = 169
y = 132
4(13) + 7 = 59 degrees. Angle one
9(13) + 4 = 121 degrees. Angle two
59 + 121 = 180 degrees.
This does check.